OpenStudy (anonymous):

five consecutive numbers having sum of 120, product of 1st and 3rd and product of 2nd and4th have a ratio of 20/21. what are they

OpenStudy (anonymous):

let numbers are x , x+1,x+2,x+3 and x+4 now as your problem says x + x+1+x+2+x+3 +x+4 =120 solve and get x........and x+4

OpenStudy (solomonzelman):

You only need the 1st sentence. lets call the MIDDLE term "x", so the 2 before would be x-2 and x-1, and the 2 after would be x+1 and x+2. Write it out, \[(x-2)+(x-1)+x+(x+1)+(x+2)=120\] solve,\[x-2+x-1+x+x+1+x+2=120\] \[5x=120\] Can you tell me the five numbers now? (hint: the middle term is the x that you are solving for)

OpenStudy (anonymous):

middle term is=120/5=24 terms are 24-2,24-1,24,24+1,24+2 note these do not satisfy second condition.

OpenStudy (solomonzelman):

Well then... but they are CONSECUTIVE and add up to 120, so these are the only possible numbers.